AS8048(1L) AS8048(1L) NAME as8048,as8041,as80c50−crossassemblersformicrocomputers SYNOPSIS as8048[-pcpu][-llistfile][-ohexfile][-d][-ssymbolfile]input as8041[-pcpu][-llistfile][-ohexfile][-d][-ssymbolfile]input etc. DESCRIPTION Theas8048,as8041commandsassembletheinputfileintoatextoutputfilerepresentingtheprogrammemoryforamicrocomputer. Options -pcpuOverridetheinstructionsetselection. ValidValuesforCpu 80C4880c4880C3580c3580C4980c4980C3980c3980C5080c5080C4080c40 80488035804980398050804080418042 -llistfile Outputafileformatedwiththememoryaddress,data,andsourceinputlines. -ohexfile Outputthememorydataimageinaformacceptedbymostpromprogrammers. -hhexfile Sameas−o. −dSavetheintermediatedatafile(seeFILE)andaborttheexecutionattheendofprocessing. -ssymbolfile Printthesymboltablevaluesandnames,oneperlineinthespecifiedfile. FILES /usr/tmp/frtXXXXXX SEEALSO FrankensteinCrossAssemblersUsersManual(filebase.doc)Appendixforas8048Frankenstein Assembler(fileas8048.doc) NOTES Thereisonlyoneinputfile. Theprogramcanselectwhichsubsetofinstructionsisaccepted.Theprogramfirstusesthename withwhichitisinvoked.Thisisscannedforasubstringwhichindicateswhichsettouse.The-p optionsoverridesthisselctionbyperformingthesamesubstringsearch.Finallytheinputfilecan selectwhichsubsettousewiththeCPUstatement. Thereshouldonlybeoneexecutablefilewithallofitsnameslinkedtoit. 1